20181223 DX ARGOS Engineering on sky

Test of Closed loop

Does the configuration change on AdSec fix the striping problem?

Poor flats

  • At 55 elevation we are only getting 191 modes in the applied flat
    • See circles in current plots
    • Edge at 10:30 is slightly rolled
    • Closed Loop with concentric circles:
      CL_with_circles.jpg
  • Actuators 573 and 582 have high forces (black with white pixels around)
  • Put these two in the position do-not-use list
    • Restart AdSec
    • Still ~190 modes
  • Move to Zenith and set flat
    • 318 modes applied
    • Highest force at act 230, -0.3
    • Slew to 47 El
    • Forces on center (act 216) goes to -0.42. OK everywhere else
    • On preset, stopAO causes flat to be applied. 173 modes
    • ARGOS closed loop see vibrations(?) on LUCI images
    • On one preset see two Ref stars on CCD47

Use FF Matrix from 20180205 (ff used when ARGOS Recs were measured)

  • cd ~aoroot/calib/adsec/current
  • mv data data.20181223
  • ln -s data data_5iter
  • Restart AdSec but FF matrix and actuator list are not consistant. FF Matrix set to null
  • Try to insert old (20180205) actuator list but still not consistant.
  • Put back correct AdSec configuration
  • Close because of sunrise
  • After closed and at Zenith set the flat with no problems and 318 modes applied

-- DougMiller - 23 Dec 2018
Topic revision: r1 - 23 Dec 2018, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback